two parts
Part 1) Find two consecutive integers whose sum is 101. Be sure and find both integers.
Part 2)Negative five subtracted from three times a number is twenty nine. Find the number

Answer by checkley79(3341)

Let x & x+2 be the 2 integers.
x+x+1=101
2x+1=101
2x=101-1
2x=100
x=100/2
x=50 ans. for the smaller integer.
50+1=51 ans. for the larger integer.
Proof:
50+51=101
101=101
------------------
3x-(-5)=29
3x+5=29
3x=29-5
3x=24
x=24/3
x=8
Proof:
3*8-(-8)=29
25+5=29
29=29